OpenStudy (anonymous):

Medal For Help :) to check my answer Simpify the expression 74 A. 343 B.16,384 C.16,807 D.2,401 ****** My answer

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

7 * 7 = 49 49 * 7 = 343 343 * 7 = 2, 401 is the answer

OpenStudy (anonymous):

Ty that is how I got I figured it too :)

OpenStudy (anonymous):

no problem good job